# Scenario Analysis with Multivariate Bayesian Machine Learning Models

### v3

@@
Contact: michael.pfarrhofer@wu.ac.at, Department of Economics, WU Vienna University of Economics and Business. We would like to thank Christiane Baumeister, Niko Hauzenberger, Philip Lane, Massimiliano Marcellino, and seminar participants at the OeNB, Swiss National Bank (SNB), and IMIM online seminar for useful comments. Disclaimer: The views expressed in this paper do not necessarily reflect those of the Oesterreichische Nationalbank or the Eurosystem.
@@ 1. Introduction
In this paper, we discuss how to conduct scenario analysis with dynamic multivariate models in macroeconomics when the functional form of the conditional mean is nonlinear and/or unknown. While related tools exist in linear and traditional nonlinear frameworks (e.g., regime switching or time-varying parameter models, see Fischer et al. (2023) for a recent example), they rely on potentially restrictive parametric assumptions and are not directly applicable in nonparametric models. This situation arises in many recent macroeconometric applications that employ Bayesian machine learning (ML) methods (see, e.g., Huber and Rossini, 2022; Clark et al., 2023; Huber et al., 2023; Hauzenberger et al., 2025b; Chernis et al., 2025; Lima et al., 2025).
@@
Conditional forecasts simulate future paths of variables under scenarios encoded as constraints on other observables, structural shocks, or both. Related techniques have been used by academics and practitioners since the 1980s in linear models (see, e.g., Doan et al., 1984), with subsequent refinements (see Waggoner and Zha, 1999; Andersson et al., 2010; Jarociński, 2010; Bańbura et al., 2015; Antolin-Diaz et al., 2021; Chernis et al., 2024; Chan et al., 2025; Crump et al., 2025). [^1] For a Bayesian decision analysis perspective and discussion of additional related literature, see West (2024). Breaking the assumption of linearity, however, complicates matters: there is no general representation of nonlinear time series as functions of shocks (see, e.g., Potter, 2000), and unknown nonlinearities make it difficult to derive multi-step ahead predictive distributions. Obtaining closed form solutions is often impossible, but in this case one can resort to Monte Carlo methods. We show how common approaches can be adapted to nonparametric models. Akin to Bańbura et al. (2015), our approach casts conditional forecasting as a nonlinear state space problem. We adopt Particle Gibbs with Ancestor Sampling (PGAS, Lindsten et al., 2014), which combines generality with computational efficiency, making it suitable for a wide range of nonlinear multivariate models.
@@
Since an IRF can be defined as the difference between conditional forecasts (Gallant et al., 1993; Koop et al., 1996; Rambachan and Shephard, 2021; Jordà and Taylor, 2025), our framework lends itself to estimating nonlinear dynamic effects (see Gonçalves et al., 2021, 2024; Kolesár and Plagborg-Møller, 2025, for related discussions) in the form of generalized IRFs (GIRFs). Compared with IRFs in linear settings (which are symmetric, shape invariant and history independent) these do not feature such potentially restrictive properties. Specifically, we consider (1) unorthogonalized GIRFs (the difference between conditional and unconditional forecasts subject to certain restrictions on observables), we explore (2) structural GIRFs to shocks identified with approaches typically used in the structural VAR (SVAR) literature, and (3) restricted GIRFs, which can be used to quantify the contributions of specific sets of transmission channels in shock propagation. [^2] A related approach derived from Antolin-Diaz et al. (2021) is discussed in Breitenlechner et al. (2022). Counterfactual experiments in this spirit relate to alternative policy rules and are subject to the Lucas (1976) critique (for a recent discussion, see McKay and Wolf, 2023). While our approach is also susceptible to these concerns, nonparametric modeling may offer flexibility and a safeguard; see also Section 2.3.
@@
The proposed approach is generally applicable in multivariate dynamic models. This means that there are many potential candidates for estimating conditional mean functions. These choices include (but are not limited to) regression trees or Gaussian process priors (see Marcellino and Pfarrhofer, 2024, for a recent review). Such nonparametric methods are more robust to misspecification that may arise from assuming tightly parameterized likelihoods, and offer the possibility of modeling genuinely nonlinear patterns in the data. For our empirical work, we use Bayesian Additive Regression Trees (BART, Chipman et al., 2010) as a specific nonparametric implementation. We pick this sum-of-trees model because tree-based approaches have proven particularly capable of producing accurate forecasts when used with time series for the US economy, with datasets structured similarly to the one we use in this paper (see, e.g., Medeiros et al., 2021; Goulet Coulombe et al., 2022; Clark et al., 2023; Goulet Coulombe, 2024). Our approach to predictive inference is developed under the assumption of an additive multivariate Gaussian error term with a time-varying covariance matrix.
@@
We illustrate our framework through three empirical applications. Our dataset comprises about $25$ quarterly macroeconomic and financial variables for the US economy ranging from the mid-$1970$s to the last quarter of $2024$. In one of our explorations, we also add some international variables for the euro area (EA) and the United Kingdom (UK). The applied work assesses and illustrates the role of nonlinearities when interest centers on CFs, and we explore asymmetries in the propagation of shocks of different types, signs and magnitudes. Specifically, we provide three empirical applications. First, inspired by Chan et al. (2025), we use a subset of the macroeconomic assumptions underlying the annual stress test conducted by the Federal Reserve System and compute CFs using constraints on observables for different scenarios, comparing predictive densities from linear and nonlinear models. Second, reflecting the growth-at-risk literature (Adrian et al., 2019), we study the unorthogonalized counterfactual implications of varying financial conditions on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock recursively (as in, e.g., Barnichon et al., 2022), and compute GIRFs to shocks of different signs and magnitudes that are allowed to propagate internationally. We then use a restricted GIRF approach to gauge the role of spillovers and spillbacks, inspired by the empirical work of Breitenlechner et al. (2022).
@@ 2.1. Predictive Simulation
where $\bm{x}_{\tau+1}=(\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$. Here, $x^{(m)}$ refers to the $m$th draw of a random variable $x$ — in most cases throughout this paper, $m$ indexes a draw from some posterior or predictive distribution when running a Markov chain Monte Carlo (MCMC) algorithm. In what follows we omit the $m$-indexing for the parameters, but stress that these computations are carried out in each sweep of the algorithm and thus account for the posterior uncertainty of parameters. For $h\geq 2$ we iterate forward, conditioning recursively on the draws for preceding horizons, by setting the predictors to $\bm{x}_{\tau+h}^{(m)}=(\bm{y}_{\tau+h-1}^{(m)\prime},\bm{y}_{\tau+h-2}^{(m)\prime},\ldots)^{\prime}$, and obtain:
@@ 2.2. Conditional Forecasts
As their name suggests, CFs imply an additional conditioning argument in (6). In this context, we denote by $\mathcal{C}_{h}$ a set that defines desired restrictions at horizon $h=1,2,\ldots,H,$ where $H$ is the maximum forecast horizon of interest, $\mathcal{C}_{1:H}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{H}\}$ collects all restrictions over the full forecast path, and the unconditional forecast results when $\mathcal{C}_{h}=\varnothing$ for all $h$. The joint distribution of interest is $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{I},\mathcal{C}_{1:H})$, and can be obtained by marginalizing $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{I},\mathcal{C}_{1:H},\bm{\Xi})$ over the parameters via MCMC sampling, as discussed previously. When there are no restrictions (i.e., for unconditional forecasts), the model can be simulated forward according to the law of motion in (1), so that any jointly realized random path over $(\tau+1):(\tau+H)$ is consistent with the model’s dynamics. In the presence of restrictions, however, the key challenge in the nonlinear context is that we cannot work directly with the joint distribution to impose them across all horizons simultaneously, as is possible in linear settings; and a decomposition like in (6) is generally not applicable.
@@
We may write these in terms of (7) by stacking both types in terms of their implications about the observables. The operator $\text{bdiag}(\bullet)$ outputs a block diagonal matrix, with the inputs arranged along the main block diagonal; we obtain $\bm{r}_{h}=[\bm{r}^{(y)}_{h};\bm{r}^{(u)}_{h}+\bm{R}^{(u)}\bm{H}\bm{\mu}_{\tau+h}]$, $\bm{R}_{h}=[\bm{R}^{(y)}_{h};\bm{R}^{(u)}_{h}\bm{H}]$ and $\bm{\Omega}_{h}=\text{bdiag}(\bm{\Omega}^{(y)}_{h},\bm{\Omega}^{(u)}_{h})$.
Particle Gibbs with Ancestor Sampling (PGAS). Under these assumptions we obtain the joint distribution of the forecast and the restrictions at horizon $h$:
@@
which encodes a nonlinear state combined with a linear measurement equation. This casts conditional forecasting and structural scenario analysis as a state space problem, similar to Bańbura et al. (2015), and allows to apply the PGAS algorithm of Lindsten et al. (2014) with a few adjustments. [^6] Our framework is also directly applicable for nowcasting problems (see, e.g., Cimadomo et al., 2022). The idea in such algorithms is to use hypothetical realizations, a set of particles $v=1,\ldots,V,$ combined with a scheme to score how plausible these particles are in light of the measurements/restrictions. A key ingredient of particle Gibbs or PGAS is that we condition on a fixed trajectory for one of the particles. This reference particle is chosen as the previous draw for the respective states in the encompassing MCMC sampler, ensuring that the Markov kernel preserves the correct invariant distribution (see Andrieu et al., 2010). The “ancestor” part of PGAS adds a resampling step for the parents of the reference trajectory, which alleviates path degeneracy by maintaining variability in the ancestral lineages, see Lindsten et al. (2014) for details. This improves mixing, especially with few particles, making the algorithm computationally attractive.
@@
While (9) places the respective restriction only at a single horizon, this conditional Gaussian distribution can be used as an “optimal proposal” within our particle algorithm, in the spirit of a forward-filtering update. This allows to generate particles that satisfy the restrictions locally without being exposed to excessive weight degeneracy (the situation in which only very few, if any, particles are compatible with the restrictions). Used within an encompassing MCMC algorithm, PGAS propagates these particles in such a way that the restrictions are enforced jointly over the entire forecast path, and we obtain draws from $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{I},\mathcal{C}_{1:H})$.
@@
    Weighting: We update the weights used in step (i) for the next horizon based on (7). Specifically, we compute the likelihood of the lineage of the respective particles in light of the restrictions. In case there are no restrictions, we trivially obtain equal weights.
@@ 2.3. Generalized Impulse Response Functions
Various types of IRFs are widely used for both academic and policy analysis. These types of dynamic effects can generally be defined as the difference between two forecasts, see Koop et al. (1996). This provides a natural link to our previous discussions on predictive distributions. Indeed, due to the recursive nonlinear structure of our model, we need to resort to a simulation-based variants of GIRFs. [^7] Simulation-based methods are often used to compute dynamic causal effects in parametric nonlinear (e.g., Baumeister and Peersman, 2013; Alessandri and Mumtaz, 2019), or nonparametric (e.g., Huber and Rossini, 2022; Clark et al., 2025; Hauzenberger et al., 2025b) models. We consider three main GIRF variants, which are nested in the expression:
@@
and differentiated by distinct conditioning assumptions. In line with the terminology in Crump et al. (2025), the first expectation refers to the “scenario” (s) forecast imposed with the restrictions $\mathcal{C}_{1:H}^{(\texttt{s})}$, and the latter is the “baseline” (b) forecast subject to $\mathcal{C}_{1:H}^{(\texttt{b})}$. Note that the conditioning on $\mathcal{I}$ here refers to the fact that these quantities are computed with draws for parameter conditioning on all available information; the expectation is taken at time $\tau$, as indicated by conditioning on $\bm{x}_{\tau+1}$. For later reference, define horizon specific GIRFs $\bm{\delta}_{\tau,h}$ based on the structure of $\bm{\Delta}_{\tau}=(\bm{\delta}_{\tau,1}^{\prime},\ldots,\bm{\delta}_{\tau,H}^{\prime})^{\prime}$. That is, the one-step ahead forecast horizon is associated with the contemporaneous impact of the restrictions (see Breitenlechner et al., 2022, for a similar timing convention).
@@
Unorthogonalized GIRF. This version is closely related to the original GIRF of Koop et al. (1996). It is obtained by restricting only observables as the scenario, which can be compared to the unconditional forecast as a baseline. Without any further restrictions, UGIRFs simply reflect a likely combination of structural shocks that drive the change in the respective observable(s), see also Crump et al. (2025). In case one is willing to impose additional structure in the model via identifying $\bm{H}$, restrictions on observables can be complemented with restrictions on shocks, see also Antolin-Diaz et al. (2021). Specifically, one may restrict a subset of the shocks to their unconditional distribution (non-driving shocks), while the unrestricted subset (driving shocks) is allowed to deviate to deliver the structural scenario in terms of the observables. In light of the assumptions about the structural form above, this implies $\bm{r}^{(u)}_{h}=\bm{0}$ and $\bm{\Omega}^{(u)}_{h}=\bm{I}$ for the non-driving shocks.
@@
The parameter $d_{0}$ is the baseline level of the shock. Since the structural errors enter linearly in our model, choices about the baseline level do not matter and we implicitly use $d_{0}=0$ in most of our subsequent discussions. Equation (11) is a special case of (10), which can be obtained by defining $\mathcal{C}_{1}^{(\texttt{s})}$ with $\bm{r}_{1}^{(u)}=(d_{0}+d)\cdot\bm{e}_{j}^{\prime}$, and $\mathcal{C}_{1}^{(\texttt{b})}$ with $\bm{r}_{1}^{(u)}=d_{0}\cdot\bm{e}_{j}^{\prime}$; $\bm{e}_{j}^{\prime}$ is the $j$th column of $\bm{I}_{n}$. For both forecasts we use $\bm{\Omega}_{1}^{(u)}=\text{diag}(1,\ldots,1,10^{-8},1,\ldots,1)$ with an approximately binding restriction in the $j$th position, $\bm{R}_{1}^{(u)}=\bm{e}_{j}$, and we have $\mathcal{C}_{h}=\varnothing$ for $h>1$. The GIRF at time $\tau$ and horizon $h$ in response to the $j$th structural shock of size $d$ is referred to as $\bm{\delta}_{j\tau,h}^{(d)}$. In Appendix A we discuss an alternative computation method which can be used without running PGAS, and Appendix B again provides an illustraion and comparison in a linear context using artificial data.
@@
Restricted GIRF. The final variant we consider combines the SGIRF with restrictions on observables, so that we may impose $\mathbb{E}(\bm{R}_{h}\bm{\delta}_{\tau,h}^{(d)})=\bm{0}_{r_{h}}$ along the desired dimensions. This approach switches off specific transmission channels of structural shocks, by partially matching the observables of the scenario forecast conditional on the shock with those of the forecasted observables in the baseline predictive distribution. Breitenlechner et al. (2022) provide a discussion of this approach in a linear context, which allows for directly restricting the IRF. In our nonlinear setting, we need to account for varying initial conditions and work with the two conditional forecasts, as we cannot directly manipulate the IRF.
@@
From an implementation perspective, we may obtain a draw $\bm{y}_{\tau+1:\tau+H}^{(\texttt{b},m)}$ from the baseline distribution $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{C}_{1}^{(\texttt{b})},\bm{x}_{\tau+1},\mathcal{I})$, where $\mathcal{C}_{1}^{(\texttt{b})}$ features the impact restrictions as in (11) plus any desired restrictions on the future sequence of non-driving shocks (this is however not a necessary requirement, as otherwise all shocks may be allowed to deviate from their unconditional distribution). For computing the scenario forecast, we then augment the shock impact restriction in (11) with a restriction on the respective observable, so that $\bm{r}_{h}^{(y)}=\bm{R}_{h}^{(y)}\bm{y}_{\tau+h}^{(\texttt{b},m)}$, to define $\mathcal{C}_{1:H}^{(\texttt{s})}$. Individual realized paths from the scenario and baseline forecast distributions are identical along the restricted dimensions. The restricted GIRF then arises as the difference in expected values as in (10), which has an expected value of zero, but this cannot strictly be enforced for each individual draw. This is because otherwise the reference trajectory in each sweep of the PGAS algorithm may not necessarily respect the restrictions.
@@
Conditional and unconditional GIRFs. One may obtain all three of these GIRFs conditional on each point in time $\tau$. In principle, one may thus consider “time-varying” dynamic effects of shocks for each period individually, or averages for different known subsets of observations (e.g., recessions vs. expansions). This time variation, however, is exclusively due to variation across initial conditions, since the functional form is nonlinear but time-invariant. Unconditional (G)IRFs can be computed in various ways (e.g., by randomizing over initial conditions and averaging, in bootstrap-type approaches); see also Kilian and Lütkepohl (2017, chapter 18), and Gonçalves et al. (2021, 2024) for definitions. In case we require estimates of unconditional versions, our preferred approach is to compute the GIRF at each point in time, and then average out this source of randomness, $\overline{\bm{\Delta}}=\sum_{\tau=1}^{T}\bm{\Delta}_{\tau}/T$.

### v4

@@
Contact: michael.pfarrhofer@wu.ac.at, Department of Economics, WU Vienna University of Economics and Business. We thank Christiane Baumeister, Niko Hauzenberger, Philip Lane, Massimiliano Marcellino, and seminar participants at the OeNB, Swiss National Bank (SNB), and IMIM seminar for useful comments. Disclaimer: The views expressed in this paper do not necessarily reflect those of the Oesterreichische Nationalbank or the Eurosystem.
@@ 1. Introduction
In this paper, we discuss how to conduct scenario analysis with dynamic multivariate models in macroeconomics when the functional form of the conditional mean is nonlinear and/or unknown. While related tools exist in linear and traditional nonlinear frameworks (e.g., regime switching or time-varying parameter models, see Fischer et al. (2023) for a recent example), they rely on potentially restrictive parametric assumptions and are not directly applicable in nonparametric models. The latter feature in many recent macroeconometric applications that employ Bayesian machine learning (ML) methods (see, e.g., Huber and Rossini, 2022; Clark et al., 2023; Huber et al., 2023; Hauzenberger et al., 2025b; Chernis et al., 2025; Lima et al., 2025).
@@
Conditional forecasts simulate future paths of variables under scenarios encoded as constraints on other observables, structural shocks, or both. Related techniques have been used by academics and practitioners since the 1980s in linear models (see, e.g., Doan et al., 1984), with subsequent refinements (see Waggoner and Zha, 1999; Andersson et al., 2010; Jarociński, 2010; Bańbura et al., 2015; Antolin-Diaz et al., 2021; Chernis et al., 2024; Chan et al., 2025). [^1] For a Bayesian decision analysis perspective and discussion of additional related literature, see West (2024). Breaking the assumption of linearity, however, complicates matters: there is no general representation of nonlinear time series as functions of shocks (see, e.g., Potter, 2000), and unknown nonlinearities make it difficult to derive multi-step ahead predictive distributions. Obtaining closed form solutions is often impossible, but in this case one can resort to Monte Carlo methods. We show how common approaches can be adapted to nonparametric settings. Akin to Bańbura et al. (2015), our approach casts conditional forecasting as a nonlinear state space problem. We use Particle Gibbs with Ancestor Sampling (PGAS, Lindsten et al., 2014), which combines generality with computational efficiency, making it suitable for a comparatively wide range of nonlinear multivariate models.
@@
An IRF can be defined as the difference between conditional forecasts (Gallant et al., 1993; Koop et al., 1996; Rambachan and Shephard, 2021; Jordà and Taylor, 2025). Our framework thus lends itself to estimating nonlinear dynamic effects (see also Gonçalves et al., 2021, 2024; Kolesár and Plagborg-Møller, 2025, for related discussions) in the form of generalized IRFs (GIRFs), and thereby contributes to improving interpretability of ML methods. Compared with IRFs in linear settings (which are symmetric, shape invariant and history independent) the GIRFs do not feature such potentially restrictive properties. Specifically, we consider (1) unorthogonalized GIRFs (the difference between conditional and unconditional forecasts subject to certain restrictions on observables), we explore (2) structural GIRFs to shocks identified with approaches typically used in the structural VAR (SVAR) literature, and (3) restricted GIRFs, which can be used to construct policy counterfactuals and quantify the contributions of specific sets of transmission channels in the propagation of shocks. [^2] A related approach derived from Antolin-Diaz et al. (2021) is discussed in Breitenlechner et al. (2022). Counterfactual experiments in this spirit relate to alternative policy rules and are subject to the Lucas (1976) critique (for a recent discussion, see McKay and Wolf, 2023). While our baseline approach is also susceptible to these concerns, there are options to improve robustness, and nonparametric modeling offers additional flexibility and a safeguard; see Section 2.3 for a more detailed discussion.
@@
The proposed approach is generally applicable in multivariate dynamic models. This means that there are many potential candidates for estimating conditional mean functions. These choices include (but are not limited to) regression trees or Gaussian process priors (see Marcellino and Pfarrhofer, 2024, for a recent review). Such nonparametric methods are more robust to misspecification that may arise from assuming tightly parameterized likelihoods, and can be used to capturing genuinely nonlinear patterns in the data. For our empirical work, we use Bayesian Additive Regression Trees (BART, Chipman et al., 2010) as a specific nonparametric implementation. We pick this sum-of-trees model because tree-based approaches have proven particularly capable of producing accurate forecasts when used with time series for the US economy, with datasets structured similarly to the one we use in this paper (see, e.g., Medeiros et al., 2021; Goulet Coulombe et al., 2022; Clark et al., 2023; Goulet Coulombe, 2024). Our approach to predictive inference is developed under the assumption of an additive multivariate Gaussian error term with a time-varying variance covariance matrix.
@@
We illustrate our framework through several empirical applications. Our dataset comprises about $25$ quarterly macroeconomic and financial variables for the US economy ranging from the mid-$1970$s to the last quarter of $2024$. In one of our explorations, we also add some international variables for the euro area (EA) and the United Kingdom (UK). The applied work assesses and illustrates the role of nonlinearities when interest centers on CFs, and we explore asymmetries in the propagation of shocks of different types, signs and magnitudes. Specifically, we provide three empirical applications. First, inspired by Chan et al. (2025), we use a subset of the macroeconomic assumptions underlying the annual stress test conducted by the Federal Reserve System and compute CFs using constraints on observables for different scenarios, comparing predictive densities from linear and nonlinear models. Second, reflecting the growth-at-risk literature (Adrian et al., 2019), we study the unorthogonalized counterfactual implications of varying financial conditions on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock recursively (as in, e.g., Barnichon et al., 2022), and compute GIRFs to shocks of different signs and magnitudes that are allowed to propagate internationally. We then use a restricted GIRF approach to gauge the role of spillovers and spillbacks, inspired by the empirical work of Breitenlechner et al. (2022).
@@ 2.1. Predictive Simulation
where $\bm{x}_{\tau+1}=(\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$. Here, $x^{(m)}$ refers to the $m$th draw of a random variable $x$ — in most cases throughout this paper, $m$ indexes a draw from a posterior or predictive distribution when running a Markov chain Monte Carlo (MCMC) algorithm. In what follows we omit the $m$-indexing for the parameters, but stress that these computations are carried out in each sweep of the algorithm and thus account for the posterior uncertainty of parameters. For $h\geq 2$ we iterate forward, conditioning recursively on the draws for preceding horizons, by setting the predictors to $\bm{x}_{\tau+h}^{(m)}=(\bm{y}_{\tau+h-1}^{(m)\prime},\bm{y}_{\tau+h-2}^{(m)\prime},\ldots)^{\prime}$, and obtain:
@@ 2.2. Conditional Forecasts
As their name suggests, CFs imply an additional conditioning argument in (6). In this context, we denote by $\mathcal{C}_{h}$ a set that defines desired restrictions at horizon $h=1,2,\ldots,H,$ where $H$ is the maximum forecast horizon of interest, $\mathcal{C}_{1:H}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{H}\}$ collects all restrictions over the full forecast path, and the unconditional forecast results when $\mathcal{C}_{h}=\varnothing$ for all $h$. The joint distribution of interest is $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{I},\mathcal{C}_{1:H})$, and can be obtained by marginalizing $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{I},\mathcal{C}_{1:H},\bm{\Xi})$ over the parameters via MCMC sampling, as discussed previously.
When there are no restrictions (i.e., for unconditional forecasts), the model can be simulated forward according to the law of motion in (1), so that any jointly realized random path over $(\tau+1):(\tau+H)$ is consistent with the model’s dynamics. In the presence of restrictions, however, the key challenge in the nonlinear context is that we cannot work directly with the joint distribution to impose them across all horizons simultaneously, as is possible in linear settings; and a decomposition like in (6) is generally not applicable.
@@
We may write these in terms of (7) by stacking both types in terms of their implications about the observables. The operator $\text{bdiag}(\bullet)$ outputs a block diagonal matrix, with the inputs arranged along the main block diagonal; we obtain $\bm{r}_{h}=[\bm{r}^{(y)}_{h};\bm{r}^{(u)}_{h}+\bm{R}^{(u)}\bm{H}\bm{\mu}_{\tau+h}]$, $\bm{R}_{h}=[\bm{R}^{(y)}_{h};\bm{R}^{(u)}_{h}\bm{H}]$ and $\bm{\Omega}_{h}=\text{bdiag}(\bm{\Omega}^{(y)}_{h},\bm{\Omega}^{(u)}_{h})$. Under these assumptions we obtain the joint distribution of the forecast and the restrictions at horizon $h$:
@@
which encodes a nonlinear state combined with a linear measurement equation. This casts conditional forecasting and structural scenario analysis as a state space problem, similar to Bańbura et al. (2015), and allows to apply the PGAS algorithm of Lindsten et al. (2014) with a few adjustments. [^6] Our framework is also applicable for nowcasting problems (see, e.g., Cimadomo et al., 2022).
Particle Gibbs with Ancestor Sampling. The idea of particle Gibbs algorithms is to use hypothetical realizations, a set of particles $v=1,\ldots,V,$ combined with a scheme to score how plausible these particles are in light of the measurements/restrictions. A key ingredient of PGAS is that we condition on a fixed trajectory for one of the particles. This reference particle is chosen as the previous draw for the respective states in the encompassing MCMC sampler, ensuring that the Markov kernel preserves the correct invariant distribution (see Andrieu et al., 2010). The “ancestor” part of PGAS adds a resampling step for the parents of the reference trajectory, which alleviates path degeneracy by maintaining variability in the ancestral lineages, see Lindsten et al. (2014) for details. This improves mixing, especially with few particles, making the algorithm computationally attractive.
@@
While (9) places the respective restriction only at a single horizon, this conditional Gaussian distribution can be used as an “optimal proposal” for particles in algorithm, in the spirit of a forward-filtering update. This allows to generate particles that satisfy the restrictions locally without being exposed to excessive weight degeneracy (the situation in which only very few, if any, particles are compatible with the restrictions). Used within an encompassing MCMC algorithm, PGAS propagates these particles in such a way that the restrictions are enforced jointly over the entire forecast path, and we obtain draws from $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{I},\mathcal{C}_{1:H})$.
@@
    Weighting: We update the weights used in step (i) for the next horizon based on (7). Specifically, we compute the likelihood of the lineage of the respective particles in light of the restrictions. In case there are no restrictions, we obtain equal weights.
@@ 2.3. Generalized Impulse Response Functions
Various types of IRFs are widely used for both academic and policy analysis. These types of dynamic effects can generally be defined as the difference between two forecasts, see Gallant et al. (1993); Koop et al. (1996). This provides a natural link to our previous discussions on predictive distributions. Indeed, due to the recursive nonlinear structure of our model, we need to resort to a simulation-based version of the GIRF. [^7] Simulation-based methods are often used to compute (generalized) impulse response functions in nonlinear parametric (e.g., Baumeister and Peersman, 2013; Alessandri and Mumtaz, 2019), or nonparametric (e.g., Huber and Rossini, 2022; Clark et al., 2025; Hauzenberger et al., 2025b) models. We consider three main variants, which are nested in the expression:
@@
and differentiated by distinct conditioning assumptions. In line with the terminology in Crump et al. (2025), the first expectation refers to the “scenario” (s) forecast imposed with the restrictions $\mathcal{C}_{1:H}^{(\texttt{s})}$, and the latter is the “baseline” (b) forecast subject to $\mathcal{C}_{1:H}^{(\texttt{b})}$. Note that the conditioning on $\mathcal{I}$ here refers to the fact that these quantities are computed with draws for parameters conditioning on all available information; the expectation is taken at time $\tau$, as indicated by conditioning on $\bm{x}_{\tau+1}$. For later reference, define horizon specific GIRFs $\bm{\delta}_{\tau,h}$ based on the structure of $\bm{\Delta}_{\tau}=(\bm{\delta}_{\tau,1}^{\prime},\ldots,\bm{\delta}_{\tau,H}^{\prime})^{\prime}$. That is, the one-step ahead forecast horizon is associated with the contemporaneous impact of the restrictions (see Breitenlechner et al., 2022, for a similar timing convention).
@@
Unorthogonalized GIRF. This version is closely related to the original GIRF of Koop et al. (1996). It is obtained by restricting only observables as the scenario $\mathcal{C}_{1:H}^{(\texttt{s})}$, which can be compared to the unconditional forecast, $\mathcal{C}_{1:H}^{(\texttt{b})}=\varnothing$, as a baseline. Without any further restrictions, UGIRFs simply reflect a likely combination of structural shocks that drive the change in the respective observable(s), see also Crump et al. (2025). In case one is willing to impose additional structure in the model via identifying structural shocks and $\bm{H}$, restrictions on observables can be complemented with restrictions on shocks (Antolin-Diaz et al., 2021). Specifically, one may restrict a subset of the shocks to their unconditional distribution (non-driving shocks), while the unrestricted subset (driving shocks) is allowed to deviate to deliver the structural scenario in terms of the observables. In light of the assumptions about the structural form above, this implies $\bm{r}^{(u)}_{h}=\bm{0}$ and $\bm{\Omega}^{(u)}_{h}=\bm{I}$ for the non-driving shocks.
@@
The parameter $d_{0}$ is the baseline level of the shock. Since the structural errors enter linearly in our model, choices about the baseline level do not matter and we implicitly use $d_{0}=0$ in most of our subsequent discussions. Equation (11) is a special case of (10), which can be obtained by defining $\mathcal{C}_{1}^{(\texttt{s})}$ with $\bm{r}_{1}^{(u)}=(d_{0}+d)\cdot\bm{e}_{j}^{\prime}$, and $\mathcal{C}_{1}^{(\texttt{b})}$ with $\bm{r}_{1}^{(u)}=d_{0}\cdot\bm{e}_{j}^{\prime}$; $\bm{e}_{j}^{\prime}$ is the $j$th column of $\bm{I}_{n}$. For both forecasts we use $\bm{\Omega}_{1}^{(u)}=\text{diag}(1,\ldots,1,10^{-8},1,\ldots,1)$ with an approximately binding restriction in the $j$th position, $\bm{R}_{1}^{(u)}=\bm{I}_{n}$, and we have $\mathcal{C}_{h}=\varnothing$ for $h>1$. The GIRF at time $\tau$ and horizon $h$ in response to the $j$th structural shock of size $d$ is referred to as $\bm{\delta}_{j\tau,h}^{(d)}$. In Appendix A we discuss an alternative computation method which can be used without running PGAS, and Appendix B again provides an illustration and comparison in a linear context using artificial data.
@@
Restricted GIRF. The final variant we consider combines the SGIRF with restrictions on observables, so that we may impose $\mathbb{E}(\bm{R}^{(y)}_{h}\bm{\delta}_{\tau,h}^{(d)})=\bm{0}_{r_{h}^{(y)}}$ along the desired dimensions. This approach switches off specific transmission channels of structural shocks, by partially matching the observables of the scenario forecast conditional on the shock with those of the forecasted observables in the baseline predictive distribution. Breitenlechner et al. (2022) provide a discussion of this approach in a linear context, which allows for directly restricting the IRF. In our nonlinear setting, we need to account for varying initial conditions and work with the two conditional forecasts, as we cannot directly manipulate the IRF.
@@
From an implementation perspective, we may obtain a draw $\bm{y}_{\tau+1:\tau+H}^{(\texttt{b},m)}$ from the baseline distribution $p(\bm{y}_{\tau+1:\tau+H}\>|\>\mathcal{C}_{1}^{(\texttt{b})},\bm{x}_{\tau+1},\mathcal{I})$, where $\mathcal{C}_{1}^{(\texttt{b})}$ features the impact restrictions as in (11) plus any desired restrictions on the future sequence of non-driving shocks (this is however not a necessary requirement, as otherwise all shocks may be allowed to deviate from their unconditional distribution; or, one can also force all of them to follow their unconditional distribution, see below). For computing the scenario forecast, we then augment the shock impact restriction in (11) with a restriction on the respective observable, so that $\bm{r}_{h}^{(y)}=\bm{R}_{h}^{(y)}\bm{y}_{\tau+h}^{(\texttt{b},m)}$, to define $\mathcal{C}_{1:H}^{(\texttt{s})}$. Individual realized paths from the scenario and baseline forecast distributions are identical along the restricted dimensions. The restricted GIRF then arises as the difference in expected values as in (10), which has an expected value of zero across MCMC iterations along the restricted dimensions, but this cannot strictly be enforced for each individual draw. This is because otherwise the reference trajectory in each sweep of the PGAS algorithm may not necessarily respect the restrictions.
@@
Conditional and unconditional GIRFs. One may obtain all three of these GIRFs conditional on each point in time $\tau$. In principle, one may thus consider “time-varying” dynamic effects of shocks for each period individually (these estimates are also referred to as “filtered” in the sense that they condition on the realized history up to a specific period, see Rambachan and Shephard, 2021). This time variation, however, is exclusively and mechanically due to variation across initial conditions, since the functional form in the class of models we consider is potentially nonlinear but time-invariant. Unconditional (G)IRFs can be computed in various ways (e.g., by randomizing over initial conditions and averaging, in bootstrap-type approaches); see also Kilian and Lütkepohl (2017, chapter 18), and Gonçalves et al. (2021, 2024) for definitions. In case we require estimates of unconditional versions, our preferred approach is to compute the GIRF at each point in time, and then average out this source of randomness, $\overline{\bm{\Delta}}=\sum_{\tau=1}^{T}\bm{\Delta}_{\tau}/T$. It is worth noting that under several assumptions, see Rambachan and Shephard (2021, Theorem 5), temporally averaged GIRFs can identify a dynamic causal effect (an “average treatment effect,” in related terminology).